The steel bed of a suspension bridge is 200 m long at 20°C. If the
extremes of temperature to which it might be exposed are -30°C
to +40°C, how much will it contract and expand?
1
Expert's answer
2019-04-01T10:35:00-0400

The difference in temperatures is equal to ∆T=70°C=70 K.

We can calculate a substantial change of length as


"ΔL=L-L0=L0(1+\alphaΔT)- L0= L0\alphaΔT (1)"

where L0=200 m, α=12×10-6 K-1


Using (1) we get: ΔL=0.168 m


Answer:

0.168 m or 17 cm